# Measuring audio prompt adherence with distribution-based embedding distances

### v1

@@ 1 Introduction
With the success of generative approaches like generative adversarial networks (GANs) and diffusion models (DM), generative AI for music has rapidly become a research topic of major interest. A variety of tools have been proposed to create musical audio, in the form of full mixes (Agostinelli et al.,, 2023; Donahue et al.,, 2023; Forsgren and Martiros,, 2022; Schneider et al.,, 2023), loops Roberts et al., (2019), individual instrument parts (Lattner and Grachten,, 2019; Grachten et al.,, 2020; Wu et al.,, 2022; Parker et al.,, 2024; Pasini et al.,, 2024), or individual instrument sounds (one-shots) Engel et al., (2017); Nistal et al., 2020b .
@@
Secondly, generated outputs are evaluated in relation to the inputs through the judgment of human subjects, often expressed as the Mean Opinion Score (MOS) Donahue et al., (2023); Parker et al., (2024). Depending on the questions to the subjects, this form of evaluation can measure a variety of aspects, including adherence of the outputs to any prompts, but due to the involvement of human subjects, is not as readily available as computational evaluation methods.
@@
Models conditioned on audio or symbolic representations of music, have been evaluated by measuring some form of reconstruction accuracy of desired characteristics, like melody, harmony, or rhythm (Wu et al., 2023a, ), or a comparison of input and output audio in terms of musical descriptors Parker et al., (2024). The former type of evaluation is most appropriate in style-transfer settings, whereas the latter is oriented towards musical accompaniment, under that assumption that stems belong to the same stem have similar distributions of musical descriptors.
@@ 2.1 Measures for audio quality
Kilgour et al., (2019) proposed the Fréchet Audio Distance as a measure of audio quality in the context of music enhancement algorithms. Rather than evaluating individual audio samples, it evaluates a set of samples as a whole. It does so by comparing the statistics of embedding representations of the set to be evaluated (the candidate set) to those of another set of samples—the reference set. Specifically, the embeddings are assumed to follow a multi-variate normal distribution, and are represented by the mean and covariance of the distribution. Kilgour et al., (2019) use the activations of the last feature layer of the VGGish model (Hershey et al.,, 2017) as embeddings. Recent experiments by Gui et al., (2023) using FAD in the context of music generation reveal that music specific embedding models, especially the CLAP joint text/audio embedding model Elizalde et al., (2023), are more effective than VGGish.
@@ 2.2 Measures for text prompt adherence
For generative models of musical audio that are conditioned on text prompts, the degree of adherence of the generated audio to the text prompt has been measured using the CLAP score. This measure is defined as the (averaged) cosine similarity between embeddings of the audio and the text in a joint text/audio embedding space defined by the CLAP model (Elizalde et al.,, 2023). The MuLan Cycle Consistency is defined similarly (Agostinelli et al.,, 2023), using the embeddings of the MuLan model Huang et al., (2022).
@@ 2.3 Other evaluation measures for conditional music generation
(Wu et al., 2023a, ) condition a generative model on symbolic melodies, and time-varying dynamics and rhythm curves. They measure output adherence to these prompts by extracting melody, dynamics and rhythm descriptors from the output audio, and compute accuracy, correlation, and F1 score, respectively.
@@ 3.2 Data collections
Tency (in-house) 19998 11.7 1255.4
Chassay (in-house) 583 10.3 28.1
Trap (in-house) 573 12.8 2.5
AT (in-house) 124 16.0 8.4
@@
The multitrack collections used in this study are listed in Table 1. The MUSDB18 (Rafii et al.,, 2019) collection is publicly available and consists of pop/rock songs. The other collections are licensed for internal use only. Of these, Tency is by far the largest with almost 20k songs in a variety of genres. The remaining datasets are smaller and consist of production music (Chassay, AT) and trap sample packs (Trap). The Trap collection consists mainly of short segments (loops) rather than full songs.
In some cases the components of drums and percussion (like kick drum, snare drum, and hi-hat) are stored as separate stems, rather than in a single stem. This accounts for the difference in steps per project between e.g. AT, and MUSDB18. Note that instruments/components typically not active throughout the entire song, so stems are often partly silent.
@@ 4.3 Results
Figure 1: Within-collection (upper three rows) and between-collection (bottom row) FAD/MMD distances of candidate sets to reference sets using different fusion methods, embedders, and projections. For between-collection distances, only the MIX fusion method is shown. The mix/stem pairs of the candidate sets are either matching (blue) or non-matching (orange).

### v2

@@ 1 Introduction
With the success of generative approaches like generative adversarial networks (GANs) and diffusion models (DM), generative AI for music has rapidly become a research topic of major interest. A variety of tools have been proposed to create musical audio, in the form of full mixes (Agostinelli et al.,, 2023; Donahue et al.,, 2023; Forsgren and Martiros,, 2022; Schneider et al.,, 2023), loops (Roberts et al.,, 2019), individual instrument parts (Lattner and Grachten,, 2019; Grachten et al.,, 2020; Wu et al.,, 2022; Parker et al.,, 2024; Pasini et al.,, 2024), or individual instrument sounds (one-shots) (Engel et al.,, 2017; Nistal et al., 2020b, ).
@@
Secondly, generated outputs are evaluated in relation to the inputs through the judgment of human subjects, often expressed as the Mean Opinion Score (MOS) (Donahue et al.,, 2023; Parker et al.,, 2024). Depending on the questions to the subjects, this form of evaluation can measure a variety of aspects, including adherence of the outputs to any prompts, but due to the involvement of human subjects, is not as readily available as computational evaluation methods.
@@
Models conditioned on audio or symbolic representations of music, have been evaluated by measuring some form of reconstruction accuracy of desired characteristics, like melody, harmony, or rhythm (Wu et al., 2023a, ), or a comparison of input and output audio in terms of musical descriptors (Parker et al.,, 2024). The former type of evaluation is most appropriate in style-transfer settings, whereas the latter is oriented towards musical accompaniment, under that assumption that stems belong to the same stem have similar distributions of musical descriptors.
@@ 2.1 Measures for audio quality
Kilgour et al., (2019) proposed the Fréchet Audio Distance as a measure of audio quality in the context of music enhancement algorithms. Rather than evaluating individual audio samples, it evaluates a set of samples as a whole. It does so by comparing the statistics of embedding representations of the set to be evaluated (the candidate set) to those of another set of samples—the reference set. Specifically, the embeddings are assumed to follow a multi-variate normal distribution, and are represented by the mean and covariance of the distribution. Kilgour et al., (2019) use the activations of the last feature layer of the VGGish model (Hershey et al.,, 2017) as embeddings. Recent experiments by Gui et al., (2023) using FAD in the context of music generation reveal that music specific embedding models, especially the CLAP joint text/audio embedding model (Elizalde et al.,, 2023), are more effective than VGGish.
@@ 2.2 Measures for text prompt adherence
For generative models of musical audio that are conditioned on text prompts, the degree of adherence of the generated audio to the text prompt has been measured using the CLAP score. This measure is defined as the (averaged) cosine similarity between embeddings of the audio and the text in a joint text/audio embedding space defined by the CLAP model (Elizalde et al.,, 2023). The MuLan Cycle Consistency is defined similarly (Agostinelli et al.,, 2023), using the embeddings of the MuLan model (Huang et al.,, 2022).
@@ 2.3 Other evaluation measures for conditional music generation
Wu et al., 2023a condition a generative model on symbolic melodies, and time-varying dynamics and rhythm curves. They measure output adherence to these prompts by extracting melody, dynamics and rhythm descriptors from the output audio, and compute accuracy, correlation, and F1 score, respectively.
@@ 3.2 Data collections
A (in-house) 19998 11.7 1255.4
B (in-house) 583 10.3 28.1
C (in-house) 124 16.0 8.4
D (in-house) 573 12.8 2.5
@@
The multitrack collections used in this study are listed in Table 1. The MUSDB18 (Rafii et al.,, 2019) collection is publicly available and consists of pop/rock songs. The other collections are licensed for internal use only. Of these, A is by far the largest with almost 20k songs in a variety of genres. The remaining collections are smaller and consist of production music (B, C) and trap sample packs (D). Collection D consists mainly of short segments (loops) rather than full songs.
In some cases the components of drums and percussion (like kick drum, snare drum, and hi-hat) are stored as separate stems, rather than in a single stem. This accounts for the difference in steps per project between e.g. C, and MUSDB18. Note that instruments/components typically not active throughout the entire song, so stems are often partly silent.
@@ 4.3 Results
Figure 1: Within-collection (upper three rows) and between-collection (bottom row) FAD/MMD distances of candidate sets to reference sets using different fusion methods, embedders, and projections. For between-collection distances, only the MIX fusion method is shown. The mix/stem pairs of the candidate sets are either matching (blue) or non-matching (orange). Asterisks denote the statistical significance of differences.